Mark has two jobs during 2018. One employer withheld and paid FICA taxes on $80,200 of Mark's salary and the other employer withheld and paid FICA taxes on $52,800 in salary paid to Mark. Calculate the amount of Mark's credit for the overpayment of Social Security taxes that he should claim on his 2018 Form 1040.
Answer #1

Social security tax rate for employees for the year 2018 =  6.2%

wage limit = 128400

Mark wages = 80200 + 52800 = 133000

Excess amount = 133000 - 128400 = 4600

Over payment = 4600*6.2% = 285.20 or 285